# Data-light Uncertainty Set Merging with Admissibility: Synthetics, Aggregation, and Test Inversion

### v2

@@ 2.1.1 Synthetic p-value
$p_{\ell}(y):=p_{\ell}(y;\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})\sim{\rm Unif}\left(0,\alpha_{\ell}\right)\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}})+{\rm Unif}\left(\alpha_{\ell},1\right)\cdot\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}}),\quad\forall y\in\mathcal{Y}.$ (2)
Intuitively, if $y\notin\mathcal{C}_{\ell,\alpha_{\ell}}$, it suggests that the underlying “oracle p-value” is likely small, so we generate the synthetic p-value from ${\rm Unif}\left(0,\alpha_{\ell}\right)$. On the other hand, if $y\in\mathcal{C}_{\ell,\alpha_{\ell}}$, it indicates that the “oracle p-value” is likely large, prompting us to generate the synthetic p-value from ${\rm Unif}\left(\alpha_{\ell},1\right)$. The following proposition demonstrates that the synthetic p-values constructed above satisfy the super-uniformity property.
@@
Suppose (1) holds. Then, we have $\mathbb{P}\{p_{\ell}(Y)\leq t\}=t+\Delta_{\ell}(t)$ for all $t\in[0,1]$ and $\ell\in[L]$, where $p_{\ell}(\cdot)$ is the mapping defined in (2) and $\Delta_{\ell}(t)=\left\{1-\mathbb{P}\left(Y\notin\mathcal{C}_{\ell,\alpha_{\ell}}\right)/{\alpha_{\ell}}\right\}\cdot\{(t-\alpha_{\ell})\operatorname{\mathds{1}}\left(t>\alpha_{\ell}\right)-(t-t\alpha_{\ell})\}/(1-\alpha_{\ell})\leq 0.$
@@
As stated in the above proposition, exact uniformity ($\mathbb{P}\{p_{\ell}(Y)\leq t\}=t$) is achieved if the uncertainty set $\mathcal{C}_{\ell,\alpha_{\ell}}$ is exact, i.e., $\mathbb{P}(Y\notin\mathcal{C}_{\ell,\alpha_{\ell}})=\alpha_{\ell}$. More generally, Proposition 1 guarantees that $p_{\ell}(Y)$ is marginally super-uniform ($\mathbb{P}\{p_{\ell}(Y)\leq t\}\leq t$). However, it is important to note that this property may not hold for a fixed candidate value $y\in\mathcal{Y}$ when considering the sampling distribution.
@@
Given an uncertainty set derived from data, the synthetic p-values defined in (2) are randomly generated. The sensitivity of the merged set to this randomness is investigated in Section F of the supplement. A naive deterministic synthetic p-value can be constructed as $\alpha_{\ell}\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}})+\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}})$. More sophisticated constructions of deterministic synthetic p-values, under certain assumptions, are detailed in Section B of the Supplement.
To compare the synthetic p-value with the “oracle p-value”, we use the following toy example to illustrate their differences in the context of merging confidence sets. Suppose we observe $X\sim\mathcal{N}(\theta^{*},1)$ and wish to construct a $(1-\alpha)$-level confidence interval for the mean parameter $\theta^{*}$. A natural choice of the “oracle p-value” for testing $H_{0}:\theta^{*}=\theta$ is $p^{\rm or}(\theta):=p^{\rm or}(\theta,X)=2\Phi(-|X-\theta|)$, and the corresponding confidence interval for $\theta^{*}$ can be written as $\mathcal{C}_{\alpha}=\{\theta:p^{\rm or}(\theta)>\alpha\}$. For the synthetic p-value generated by $p(\theta)\sim{\rm Unif}\left(0,\alpha\right)\cdot\operatorname{\mathds{1}}(\theta\notin\mathcal{C}_{\alpha})+{\rm Unif}\left(\alpha,1\right)\cdot\operatorname{\mathds{1}}(\theta\in\mathcal{C}_{\alpha})$, its comparison with $p^{\rm or}(\theta)$ is provided in Figure 1.
[…]
Figure 1: Comparison of oracle p-value (blue solid) and synthetic p-value (red dot). Left: null comparison with $\Delta(t)=F^{\rm or}(t)-F(t)$, where $F^{\rm or}(\cdot)$ and $F(\cdot)$ are CDFs of oracle and synthetic p-values, respectively. Middle & right: non-null comparisons with $\theta=1$ and $\theta^{*}=0$.
The left panel compares the CDFs of $p^{\rm or}(\theta)$ and $p(\theta)$ when $\theta=\theta^{*}$, showing that $p^{\rm or}(\theta)$ and $p(\theta)$ share the same uniform distribution in this case. The middle and right panels respectively compare the CDFs and density functions of $p^{\rm or}(\theta)$ and $p(\theta)$ when $\theta=1$ and $\theta^{*}=0$. We observe that $p^{\rm or}(\theta)$ is stochastically smaller than $p(\theta)$, indicating that $p^{\rm or}(\theta)$ is more powerful for testing $H_{0}:\theta^{*}=\theta$, while $p(\theta)$ offers a good approximation.
2.1.2 Synthetic e-value
Recently, e-value has gained popularity for hypothesis testing due to its favorable properties (Vovk and Wang,, 2021; Shafer,, 2021; Grünwald et al.,, 2020; Wang and Ramdas,, 2022). An e-value is a non-negative random variable with an expected value no greater than $1$ under the null hypothesis. Again, the true underlying e-value for each study, which we refer to as “oracle e-value”, is not available under our framework. Thus, for each uncertainty set $\mathcal{C}_{\ell,\alpha_{\ell}}$ with control level $\alpha_{\ell}$, $\ell\in[L]$, we propose to generate a synthetic e-value as follows:
$e_{\ell}(y):=e(y;\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})=\alpha_{\ell}^{-1}\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell,\alpha_{\ell}}),\quad\forall y\in\mathcal{Y},$ (3)
where the corresponding function $e_{\ell}(\cdot)$ is referred to as an e-function. The expectation of $e_{\ell}(Y)$ is analyzed in the proposition below.
@@ 2.1.2 Synthetic e-value
Remark 3.
The expectation in Proposition 2 is taken with respect to the randomness in both $Y$ and $\mathcal{C}_{\ell,\alpha_{\ell}}$. Notably, existing work primarily focuses on constructing e-values based on testing results with well-defined null hypotheses (e.g., Ren and Barber,, 2024; Bashari et al.,, 2024). In contrast, our framework differs considerably because establishing a well-defined null hypothesis in the current context is challenging due to the potential randomness of $Y$.
2.2 Aggregation of Synthetic Statistics
In this section, we provide a detailed discussion on how to aggregate synthetic p-values and e-values. The aggregated statistics will then be transformed back into a merged uncertainty set in the next step. Define $\mathbf{p}(Y)=\{p_{1}(Y),\dots,p_{L}(Y)\}\in[0,1]^{L}$ and $\mathbf{e}(Y)=\{e_{1}(Y),\dots,e_{L}(Y)\}\in[0,\infty)^{L}$, where $\mathbf{p}(\cdot)$ and $\mathbf{e}(\cdot)$ are mappings specified in Algorithm 1.
2.2.1 Aggregation under independence
We first consider independent case. Let $\mathbf{p}=(p_{1},\ldots,p_{L})$ be a vector of independent synthetic p-values. If each $p_{\ell}$ is a conventional p-value, then as pointed out in Vovk et al., (2022), we can define an aggregation function $G_{p}:[0,1]^{L}\mapsto[0,1]$ through the corresponding rejection regions. More precisely, given any increasing collection of Borel lower sets $\{R_{\alpha}\subseteq[0,1]^{L}:\alpha\in(0,1)\}$, if $\mathbb{P}(\mathbf{p}\in R_{\alpha})\leq\alpha$ for any $\alpha\in(0,1)$ under the null hypothesis, then $G_{p}(\mathbf{p})=\inf\{\alpha\in(0,1):\mathbf{p}\in R_{\alpha}\}$ defines a valid p-value aggregation function. Synthetic p-values can be aggregated using the same idea. Specifically, we consider the rejection regions of the following form:
@@ 2.2.1 Aggregation under independence
where $S_{\ell}:[0,1]\mapsto\mathbb{R}$ is decreasing and $c_{1-\alpha}(\{S_{\ell}\}_{\ell\in[L]})={\rm Quantile}(1-\alpha;\sum_{\ell=1}^{L}S_{\ell}(U_{\ell}))$ with $U_{\ell}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}{\rm Unif}[0,1]$. Notably, (4) encompasses some of the most widely used aggregation methods for the conventional p-values. For example, by setting $S_{\ell}(t)=-2\log t$, we obtain Fisher’s aggregation function $G_{p}(\mathbf{p})=1-F_{\chi^{2}_{2L}}(-2\sum_{\ell=1}^{L}\log p_{\ell})$, where $F_{\chi^{2}_{2L}}$ denotes the CDF of a centered $\chi^{2}$-random variable with $2L$ degrees of freedom (Fisher,, 1948); by taking $S_{\ell}(t)=-\lambda_{\ell}\cdot\Phi^{-1}(t)$, where $\lambda_{\ell}$’s are some positive constants, we obtain the Lipták’s method $G_{p}(\mathbf{p})=\Phi\Big{\{}\sum_{\ell=1}^{L}\lambda_{\ell}\cdot\Phi^{-1}(p_{\ell})/\sqrt{\sum_{\ell=1}^{L}\lambda_{\ell}^{2}}\Big{\}}$ (Lipták,, 1958). For detailed comparisons of various aggregation methods under independence, see, for example, Heard and Rubin-Delanchy, (2018).
@@
Suppose (1) holds. Let $\bar{p}(Y):=G_{p}\{\mathbf{p}(Y)\}=\inf\{\alpha\in(0,1):\mathbf{p}(Y)\in R_{\alpha}\}$, where $R_{\alpha}$ is defined in (4). If the entries of $\mathbf{p}(Y)$ are mutually independent, then $\mathbb{P}\{\bar{p}(Y)\leq t\}\leq t$ for all $t\in[0,1]$.
We next turn to the aggregation of independent synthetic e-values. The most popular method for aggregating independent conventional e-values is through multiplication or averaging (Vovk and Wang,, 2021). Similarly, synthetic e-values can be aggregated by
$\bar{e}_{k}:=G_{e}(\mathbf{e};k)=\binom{L}{k}^{-1}\sum_{\mathcal{I}_{k}\in\mathcal{B}_{k}}\prod_{\ell\in\mathcal{I}_{k}}e_{\ell},$ (5)
where $\mathbf{e}=(e_{1},\ldots,e_{L})$ is a vector of independent synthetic e-values and $\mathcal{B}_{k}$ is the set of all $k$-element subsets of $[L]$. We abbreviate $\bar{e}_{k}$ and $G_{e}(\cdot;k)$ as $\bar{e}$ and $G_{e}(\cdot)$, respectively, whenever $k$ is fixed and does not influence the calculations. The validity of this aggregation method is established by the following proposition.
@@
Suppose (1) holds. Let $\bar{e}_{k}(Y)=G_{e}\{\mathbf{e}(Y);k\}$ be defined as in (5). If the entries of $\mathbf{e}(Y)$ are mutually independent, then $\mathbb{E}\{\bar{e}_{k}(Y)\}\leq 1$ for any pre-determined $k\in[L]$.
When $Y$ is a fixed but unknown parameter and each of the $L$ studies independently collects data to construct $\mathcal{C}_{\ell,\alpha_{\ell}}$, the entries in $\mathbf{p}(Y)$ and $\mathbf{e}(Y)$ are independent, making the aggregation methods discussed in this section appropriate. However, if $Y$ is random, the entries in $\mathbf{p}(Y)$ and $\mathbf{e}(Y)$ are generally not independent, even if each study independently collects data to construct $\mathcal{C}_{\ell,\alpha_{\ell}}$. This is because the randomness of $Y$ introduces dependence among the entries via their shared dependency on $Y$. Thus, it is crucial to account for this to ensure the validity of the aggregation method. The aggregation approach under dependence will be discussed next.
2.2.2 Aggregation under dependence
We now explore methods for aggregating synthetic statistics under arbitrary dependence. Note that, more sophisticated aggregation methods can be employed if specific dependence structures are assumed as discussed in Section A.2 of the supplement. Let $\mathbf{p}=(p_{1},\ldots,p_{L})$ be a vector of possibly dependent synthetic p-values. Consider a family of regions defined as follows:
where $\lambda_{\ell}$’s are non-negative numbers satisfying $\sum_{\ell=1}^{L}\lambda_{\ell}=1$, and $f_{\ell}$’s are p-to-e calibrators. Here, p-to-e calibrator is a decreasing function $f:[0,\infty)\mapsto[0,\infty]$ such that $\|f\|_{L_{1}([0,1])}\leq 1$ (Vovk and Wang,, 2021; Gasparin et al.,, 2024). This form of p-value aggregation is proposed in Vovk et al., (2022). If $p_{\ell}$’s are conventional p-values, $G_{p}(\mathbf{p})=\inf\{\alpha\in(0,1):\mathbf{p}\in R_{\alpha}\}$ with $R_{\alpha}$ defined in (6) encompasses some popular aggregation methods. For instance, if $f_{\ell}(p)=2-2p$ and $\lambda_{\ell}=1/L$ for all $\ell\in[L]$, we get the arithmetic mean aggregation function $G_{p}(\mathbf{p})=2\sum_{\ell=1}^{L}p_{\ell}/L$; for a pre-determined $k\in[L]$, if $f_{\ell}(p)=L/k\cdot\operatorname{\mathds{1}}\{p\in(0,k/L)\}+\infty\operatorname{\mathds{1}}(p=0)$ and $\lambda_{\ell}=1/L$ for all $\ell\in[L]$, we obtain the Rüger’s method $G_{p}(\mathbf{p})=L/k\cdot p_{(k)}$.
Analogous to Proposition 3, the following proposition establishes the super-uniformity of the aggregated synthetic p-values.
Proposition 5.
Suppose (1) holds. Let $\bar{p}(Y):=G_{p}\{\mathbf{p}(Y)\}=\inf\{\alpha\in(0,1):\mathbf{p}(Y)\in R_{\alpha}\}$, where $R_{\alpha}$ is defined in (6). Then, we have $\mathbb{P}\{\bar{p}(Y)\leq t\}\leq t$ for all $t\in[0,1]$.
For synthetic e-values under dependence, aggregation can be achieved through convex combination. The theoretical guarantee is provided in the next proposition.
Proposition 6.
Suppose (1) holds. Let $\bar{e}(Y):=\bar{e}(Y;\lambda)=\sum_{\ell=1}^{L}\lambda_{\ell}e_{\ell}(Y)$, where $\lambda_{\ell}\geq 0$ for all $\ell$ and $\sum_{\ell=1}^{L}\lambda_{\ell}=1$. Then, we have $\mathbb{E}\{\bar{e}(Y)\}\leq 1$.
2.3 Test Inversion
As a final step, we transform the aggregated synthetic statistics back into an uncertainty set. It is well known that there is a correspondence between hypothesis testing and set estimation. In fact, a standard method for constructing confidence sets is through test inversion (Casella and Berger,, 2024). We state this classical result in the following proposition.
Proposition 7.
Suppose we have a $\alpha$-level test for the null hypothesis $H_{0}(\theta):\theta^{*}=\theta$ for each $\theta\in\Theta$. If we define $\mathcal{C}_{\alpha}=\{\theta:H_{0}(\theta)\ \text{is not rejected}\}$, then $\mathcal{C}_{\alpha}$ is a valid $1-\alpha$ confidence set.
The inversion principle similarly applies to constructing general uncertainty sets, as formally stated later in Proposition 8. Based on this observation, we perform the test inversion step of Algorithm 1 based on the aggregated p-values or e-values. Specifically, for any pre-specified $\alpha\in(0,1)$, the final merged set is constructed via
$\bar{\mathcal{C}}_{\alpha}=\{y:\bar{p}(y)>\alpha\}\text{ or }\bar{\mathcal{C}}_{\alpha}=\{y:\bar{e}(y)<\tau/\alpha\},$
where $\bar{p}(y)$ and $\bar{e}(y)$ are the aggregated statistics studied in Section 2.2.
Remark 4.
Consider the scenario where $\alpha_{\ell}=\alpha$ for all $\ell\in[L]$. If $\tau=1/2$ and $\bar{e}(y)=\frac{1}{L}\sum_{\ell=1}^{L}e_{\ell}(y)$, then the merged set $\bar{\mathcal{C}}_{\alpha}=\{y:\bar{e}(y)<\tau/\alpha\}$ reduces to the majority voting set $\bar{\mathcal{C}}_{\alpha}^{\tt MV}=\big{\{}y\in\mathcal{Y}:\frac{1}{L}\sum_{\ell=1}^{L}\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell,\alpha_{\ell}})\geq\frac{1}{2}\big{\}}$ proposed in Gasparin and Ramdas, (2024).
Remark 5.
Randomization techniques for power improvement, similar to those presented in Gasparin et al., (2024); Gasparin and Ramdas, (2024), can also be applied to our construction of synthetic statistics. Specifically, we can replace (6) with
$R^{\rm U}_{\alpha}=\left\{\mathbf{p}\in[0,1]^{L}:\sum_{\ell=1}^{L}\lambda_{\ell}\cdot f_{\ell}\left(\frac{p_{\ell}}{\alpha}\right)\geq U\right\},$
and $\bar{e}(Y;\lambda)$ with $\bar{e}(Y;\lambda)/U$, where $U$ is an independently generated $\rm{Unif}(0,1)$ random variable. By employing such randomization, we may obtain smaller merged sets with valid coverage rates. The validity of these procedures follows from Theorem 3.10 in Gasparin et al., (2024) and Theorem 1.2 in Ramdas and Manole, (2023).

### v3

@@ 3.1 Synthetic p-values
$p_{\ell}(y):=p_{\ell}(y;\mathcal{C}_{\ell},\alpha_{\ell})\sim\operatorname{Unif}\left(0,\alpha_{\ell}\right)\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell})+\operatorname{Unif}\left(\alpha_{\ell},1\right)\cdot\operatorname{\mathds{1}}(y\in\mathcal{C}_{\ell}),\quad\forall y\in\mathcal{Y}.$ (4)
The construction is intuitive: if a point $y$ is outside the set $\mathcal{C}_{\ell}$, the oracle p-value was likely small, so we draw a value from $\operatorname{Unif}(0,\alpha_{\ell})$. Conversely, if $y$ is inside the set, the oracle p-value was likely large, and we draw from $\operatorname{Unif}(\alpha_{\ell},1)$. The following proposition confirms that this construction yields a valid, super-uniform p-value.
@@
Suppose (1) holds. Then, we have $\mathbb{P}\{p_{\ell}(Y)\leq t\}=t+\Delta_{\ell}(t)$ for all $t\in[0,1]$ and $\ell\in[L]$, where $p_{\ell}(\cdot)$ is the mapping defined in (4) and $\Delta_{\ell}(t)=\left\{1-\mathbb{P}\left(Y\notin\mathcal{C}_{\ell}\right)/{\alpha_{\ell}}\right\}\cdot\{(t-\alpha_{\ell})\operatorname{\mathds{1}}\left(t>\alpha_{\ell}\right)-(t-t\alpha_{\ell})\}/(1-\alpha_{\ell})\leq 0.$
@@
As stated in the above proposition, exact uniformity ($\mathbb{P}\{p_{\ell}(Y)\leq t\}=t$) is achieved if the uncertainty set $\mathcal{C}_{\ell}$ is exact, i.e., $\mathbb{P}(Y\notin\mathcal{C}_{\ell})=\alpha_{\ell}$.
@@
Given an uncertainty set derived from data, the synthetic p-values defined in (4) are randomly generated. The sensitivity of the merged set to this randomness is investigated in Section E of the Supplement.
3.2 Aggregating Synthetic p-values and Test Inversion
Having constructed synthetic p-values, the next step is to combine them into a single aggregated p-value, $\bar{p}(y)$. The choice of aggregation method depends on the dependence structure of the p-values.
We first address the general case of arbitrary dependence. Let $\mathbf{p}=(p_{1},\ldots,p_{L})$ be a vector of synthetic p-values. If each $p_{\ell}$ is a conventional p-value, then as pointed out in Vovk et al., (2022), we can define an aggregation function $G_{p}:[0,1]^{L}\mapsto[0,1]$ through the corresponding rejection regions. More precisely, given any increasing collection of Borel lower sets $\{R_{\alpha}\subseteq[0,1]^{L}:\alpha\in(0,1)\}$, if $\mathbb{P}(\mathbf{p}\in R_{\alpha})\leq\alpha$ for any $\alpha\in(0,1)$ under the null hypothesis, then $G_{p}(\mathbf{p})=\inf\{\alpha\in(0,1):\mathbf{p}\in R_{\alpha}\}$ defines a valid p-value aggregation function. Synthetic p-values can be aggregated using the same idea.
Consider a family of regions defined as follows:
where $\lambda_{\ell}$’s are non-negative numbers satisfying $\sum_{\ell=1}^{L}\lambda_{\ell}=1$, and $f_{\ell}$’s are p-to-e calibrators. Here, p-to-e calibrator is a decreasing function $f:[0,\infty)\mapsto[0,\infty]$ such that $\|f\|_{L_{1}([0,1])}\leq 1$ (Vovk and Wang,, 2021; Gasparin et al.,, 2024). This form of p-value aggregation is proposed in Vovk et al., (2022). If $p_{\ell}$’s are conventional p-values, $G_{p}(\mathbf{p})=\inf\{\alpha\in(0,1):\mathbf{p}\in R_{\alpha}\}$ with $R_{\alpha}$ defined in (5) encompasses some popular aggregation methods. For instance, if $f_{\ell}(p)=2-2p$ and $\lambda_{\ell}=1/L$ for all $\ell\in[L]$, we get the arithmetic mean aggregation function $G_{p}(\mathbf{p})=2\sum_{\ell=1}^{L}p_{\ell}/L$; for a pre-determined $k\in[L]$, if $f_{\ell}(p)=L/k\cdot\operatorname{\mathds{1}}\{p\in(0,k/L)\}+\infty\operatorname{\mathds{1}}(p=0)$ and $\lambda_{\ell}=1/L$ for all $\ell\in[L]$, we obtain the Rüger’s method $G_{p}(\mathbf{p})=L/k\cdot p_{(k)}$. The following proposition establishes the super-uniformity of the aggregated synthetic p-values.
@@ 3.2 Aggregating Synthetic p-values and Test Inversion
Suppose $\mathbb{P}\{p_{\ell}(Y)\leq t\}\leq t$ for all $\ell\in[L]$. Let $\bar{p}(Y):=G_{p}\{\mathbf{p}(Y)\}=\inf\{\alpha\in(0,1):\mathbf{p}(Y)\in R_{\alpha}\}$, where $R_{\alpha}$ is defined in (5). Then, we have $\mathbb{P}\{\bar{p}(Y)\leq t\}\leq t$ for all $t\in[0,1]$.
When the synthetic p-values are known to be independent, more powerful aggregation methods are available. Specifically, we consider the rejection regions of the following form:
@@
where $S_{\ell}:[0,1]\mapsto\mathbb{R}$ is decreasing and $c_{1-\alpha}(\{S_{\ell}\}_{\ell\in[L]})=\mathrm{Quantile}(1-\alpha;\sum_{\ell=1}^{L}S_{\ell}(U_{\ell}))$ with $U_{\ell}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\operatorname{Unif}(0,1)$. Notably, (6) encompasses some of the most widely used aggregation methods for the conventional p-values. For example, by setting $S_{\ell}(t)=-2\log t$, we obtain Fisher’s aggregation function $G_{p}(\mathbf{p})=1-F_{\chi^{2}_{2L}}(-2\sum_{\ell=1}^{L}\log p_{\ell})$, where $F_{\chi^{2}_{2L}}$ denotes the CDF of a centered $\chi^{2}$-random variable with $2L$ degrees of freedom (Fisher,, 1948); by taking $S_{\ell}(t)=-\lambda_{\ell}\cdot\Phi^{-1}(t)$, where $\lambda_{\ell}$’s are some positive constants, we obtain the Lipták’s method $G_{p}(\mathbf{p})=\Phi\Big\{\sum_{\ell=1}^{L}\lambda_{\ell}\cdot\Phi^{-1}(p_{\ell})/\sqrt{\sum_{\ell=1}^{L}\lambda_{\ell}^{2}}\Big\}$ (Lipták,, 1958). For detailed comparisons of various aggregation methods under independence, see, for example, Heard and Rubin-Delanchy, (2018).
@@
Suppose $\mathbb{P}\{p_{\ell}(Y)\leq t\}\leq t$ for all $\ell\in[L]$. Let $\bar{p}(Y):=G_{p}\{\mathbf{p}(Y)\}=\inf\{\alpha\in(0,1):\mathbf{p}(Y)\in R_{\alpha}\}$, where $R_{\alpha}$ is defined in (6). If the entries of $\mathbf{p}(Y)$ are mutually independent, then $\mathbb{P}\{\bar{p}(Y)\leq t\}\leq t$ for all $t\in[0,1]$.
The final step of the randomized procedure is to convert the aggregated p-value, $\bar{p}(y)$, back into a merged set via test inversion. A p-value controls the Type I error by rejecting the null hypothesis when $p\leq\alpha$. Inverting this decision rule directly yields the corresponding uncertainty set.
@@
Applying Proposition 8, the final merged set is constructed as
$\bar{\mathcal{C}}_{\alpha}=\{y\in\mathcal{Y}:\bar{p}(y)>\alpha\},$
where $\bar{p}(y)$ is a valid aggregated p-value from either (5) or (6).
The validity of the above $\bar{\mathcal{C}}_{\alpha}$ is a direct consequence of the validity of its constituent steps, and is formalized in Theorem 1 in the previous section.